OpenStudy (anonymous):

a^2-2a+1/6 divided by a^2-1

OpenStudy (saifoo.khan):

Please use brackets.

OpenStudy (anonymous):

its kinda obvious in this case lol

OpenStudy (anonymous):

(a^2-2a+1) / 6 / [ a^2-1]

OpenStudy (anonymous):

happy lol, u can see the cancellation

OpenStudy (anonymous):

\[= \frac{(a-1)^2}{6} \times \frac{1}{(a-1)(a+1)} \]\[= \frac{(a-1)}{6(a+1)}\]